40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Apr 29 22:53:06 2018, gautam, Update, General, DARM actuation estimate darmEst.pdf
    Reply  Tue May 1 19:37:50 2018, gautam, Update, General, DARM actuation estimate darmEst.pdfdarmEst_time.pdf
       Reply  Mon May 7 16:23:06 2018, gautam, Update, General, DARM actuation estimate darmProj.pdf37.pngMCalign_20180507.png
          Reply  Mon May 7 20:01:14 2018, Rorpheus, Update, General, Use anti-dewhitening + show CARMA/DARMA 
             Reply  Thu May 10 14:13:22 2018, gautam, Update, General, More refinement of DARM control signal projection darmProj.pdfdarmOLTF.pdfDWcomparison.pdf
                Reply  Fri May 11 13:58:42 2018, rana, Update, General, More refinement of DARM control signal projection 
                   Reply  Fri May 11 19:02:52 2018, gautam, Update, General, More refinement of DARM control signal projection darmProj.pdfdarmOLTF.pdfDWcomparison.pdfDARMsuppression.pdfControlLoop.pdf
                      Reply  Sat May 12 10:02:03 2018, rana, Update, General, More refinement of DARM control signal projection 
                         Reply  Sun May 20 17:43:01 2018, rana, Update, Electronics, How to choose resistors 
                      Reply  Mon May 14 18:58:32 2018, Kevin, Update, PonderSqueeze, Squeezing with no SRM homodyne_heatmap.pdfdisplacement_noise.pdfnoise_budget.pdf
                         Reply  Tue Oct 2 23:57:16 2018, gautam, Update, PonderSqueeze, Squeezing scenarios PonderSqueeze_NB_LORIN.pdf
                   Reply  Tue May 15 21:56:57 2018, gautam, Update, General, Stack measurement setup decommissioned 
                      Reply  Thu May 17 09:14:38 2018, Steve, Update, General, Stack measurement setup decommissioned Load_Cell_Measurement_Set_Up.jpgETMX_stack_up_down.png
Message ID: 14225     Entry time: Tue Oct 2 23:57:16 2018     In reply to: 13841
Author: gautam 
Type: Update 
Category: PonderSqueeze 
Subject: Squeezing scenarios 

[kevin, gautam]

We have been working on double checking the noise budget calculations. We wanted to evaluate the amount of squeezing for a few different scenarios that vary in cost and time. Here are the findings:

Squeezing scenarios

Sqz [dBvac] fmin [Hz] PPRM [W] PBS [W] TPRM [%] TSRM [%]
-0.41 215 0.8 40 5.637 9.903
-0.58 230 1.7 80 5.637 9.903
-1.05 250 1.7 150 1 17
-2.26 340 10 900 1 17

All calculations done with

  • 4.5kohm series resistance on ETMs, 15kohms on ITMs, 25kohm on slow path on all four TMs.
  • Detuning of SRC = -0.01 deg.
  • Homodyne angle = 89.5 deg.
  • Homodyne QE = 0.9. 
  • Arm losses is 20ppm RT.
  • LO beam assumed to be extracted from PR2 transmission, and is ~20ppm of circulating power in PRC.

Scenarios:

  1. Existing setup, new RC folding mirrors for PRG of ~45.
  2. Existing setup, send Innolight (Edwin) for repair (= diode replacement?) and hope we get 1.7 W on back of PRM.
  3. Repair Innolight, new PRM and SRM, former for higher PRG, latter for higher DARM pole.
  4. Same as #3, but with 10 W input power on back of PRM (i.e. assuming we get a fiber amp).

Remarks:

  • The errors on the small dB numbers is large - 1% change in model parameters (e.g. arm losses, PRG, coil driver noise etc) can mean no observable squeezing. 
  • Actually, this entire discussion is moot unless we can get the RIN of the light incident on the PRM lower than the current level (estimated from MC2 transmission, filtered by CARM pole and ARM zero) by a factor of 60dB.
    • This is because even if we have 1mW contrast defect light leaking through the OMC, the beating of this field (in the amplitude quadrature) with the 20mW LO RIN (also almost entirely in the amplitude quad) yields significant noise contribution at 100 Hz (see Attachment #1).
    • Actually, we could have much more contrast defect leakage, as we have not accounted for asymmetries like arm loss imbalance.
    • So we need an ISS that has 60dB of gain at 100 Hz. 
    • The requirement on LO RIN is consistent with Eq 12 of this paper.
  • There is probably room to optimize SRC detuning and homodyne angle for each of these scenarios - for now, we just took the optimized combo for scenario #1 for evaluating all four scenarios.
  • OMC displacement noise seems to only be at the level of 1e-22 m/rtHz, assuming that the detuning for s-pol and p-pol is ~30 kHz if we were to lock at the middle of the two resonances
    • This assumes 0.02 deg difference in amplitude reflectivity b/w polarizations per optic, other parameters taken from aLIGO OMC design numbers.
    • We took OMC displacement noise from here.

Main unbudgeted noises:

  • Scattered light.
  • Angular control noise reinjection (not sure about the RP angular dynamics for the higher power yet).
  • Shot noise due to vacuum leaking from sym port (= DC contrast defect), but we expect this to not be significant at the level of the other noises in Atm #1.
  • Osc amp / phase.
  • AUX DoF cross coupling into DARM readout.
  • Laser frequency noise (although we should be immune to this because of our homodyne angle choice).

Threat matrix has been updated.

Attachment 1: PonderSqueeze_NB_LORIN.pdf  184 kB  Uploaded Wed Oct 3 01:22:32 2018  | Hide | Hide all
PonderSqueeze_NB_LORIN.pdf
ELOG V3.1.3-